• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh Subedi

    Log entry time 08:08:05 on March18,2005

    Entry number 141814

    keyword=Shift Summary

    00:00 Shift started with no beam.
    00:12 MCC made a call saying beam will be back in a minute.

    All production runs in slanted carbon at 40uA beam.
    00:14-01:04 Run 2754(200K)
    01:04-01:52 Run 2755(200K)
    01:53-02:46 Run 2756(200K)
    02:47-03:40 Run 2757(200K)

    Observed randomly fluctuating deadtime regardless of beam or no beam.

    03:20 Seening many beam trips (about 30 trips per hour),called MCC to decrease the current if the trip is due to current-related problems, they replied that the trip is current-independent and actually the RF-related problem and decreasing current doesn't help it.

    03:42-04:29 Run 2758(200K)
    04:30-05:52 Run 2759(200K)
    In online display of run 2759 tdc values in BB (aux- delta E and E-planes only, ND was ok) were very low.

    05:26-05:52 Run 2760(95k)

    Eli was curious about the halog entry 141799 and 141800 made about the low tdc values of BB and telephoned me at 05:45. As per his suggestion run 2760 was ended early to reset ROC6 and to do "startcoda" so as to see if it was the reason for low tdc values in BB
    After finishing this, analyzed the new run 2761, it was OK

    Interestingly, the run 2760 had already good BB tdc spectra.

    Problematic runs (low tdc value BB): 2722,2752,2759.


    I counted the number of beam trips from 5 to 6 o'clock , it was 29.Same was the case in 2 to 3 o'clock too.

    05:56-06:42 Run 2761 (200k)
    06:43-07:40 Run 2762 (200k)

    07:10 SRC HV gui in Hall A General Tools crashed. The BB/ND-scaler display is OK, HV is still surviving.


    07:41- Run 2763, passed to day shift.